OpenStudy (anonymous):

Five spoons and two glasses weigh as much as 3 plates. One plate weighs as much as 1 spoon and 1 glass. How many spoons weigh as much as 1 glass?

OpenStudy (anonymous):

set variables, s = spoons g = glass p = plates 5s + 2g = 3p 1p = 1s +1g #s = 1g

OpenStudy (anonymous):

looking at the 2nd equation above...I would multiply each side by 3 3p = 3s + 3g i did that so that I can set equation 1 and 2 to equal each other

OpenStudy (anonymous):

5s + 2g = 3s + 3g 5s - 3s = 3g - 2g 2s = 1g so your answer is 2 spoons weigh as much as 1 glass

OpenStudy (anonymous):

Thank you so much!